accelerator complex status
Download
Skip this Video
Download Presentation
Accelerator Complex Status

Loading in 2 Seconds...

play fullscreen
1 / 7

Accelerator Complex Status - PowerPoint PPT Presentation


  • 108 Views
  • Uploaded on

Accelerator Complex Status. P. Collier. Linac2, Booster and PS Startup on-time, according to the schedule. Only minor problems Rapidly set-up the major beams for Physics users and LHC. CT transfer setup to SPS (MTE will come later)

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about 'Accelerator Complex Status' - yair


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
slide2
Linac2, Booster and PS
  • Startup on-time, according to the schedule.
  • Only minor problems
  • Rapidly set-up the major beams for Physics users and LHC.
  • CT transfer setup to SPS (MTE will come later)
  • Initially limited to low intensity in the PS due to work on the POPS

X

  • SPS
  • Originally had a 1-week delay due to additional damaged cables in LSS2 discovered during the cable replacement campaign
  • Careful rescheduling during the machine-checkout allowed all activities to be completed more quickly
  • First beam could be taken on 23rd April as originally foreseen.
slide3
Problems During the Start-up Period
  • Isolde :
  • Problem with the motorization of the GPS front end. Orphan equipment. Managed to find a spare card. Equipment now been taken over by EN-STI.
  • Rex run delayed due to a vacuum problem in the linac.
  • AD:
  • Target replacement revealed a problem with the target position – 6cm longitudinal misalignment. Was there during the whole of the 2008 run. Repaired in situ and re-aligned correctly.
  • SPS:
  • ZS (Electrostatic septum) replacement just before beam start. Had to be re-opened because of an alignment problem.
  • Inter-turn short in a main dipole developed during early operation. Magnet changed.
  • PSB:
  • Injection Septum developed a water leak into the vacuum.
    • Had run without problem for 30 years!
    • Spare magnets available – but as old as the original. Difficulties getting drawings together and tooling.
    • A few experts on the equipment still around to help.
    • Total time for repair 1 week.
slide4
radiation levels

damaged zone

Not only radiation but also risk of contamination

(overalls/gloves obligatory for work on “inside” components)

Courtesy: J. Borburgh, V. Mertens TE/ABT

slide5
Latest News
  • CNGS:
  • Set-up of the beam on the target and secondary beam optimization was delayed by the 1 week stop in the PSB.
  • However, very smooth startup once beam was available.
  • On Monday 1st June (start of CNGS physics) had 3xCNGS cycles with ~2x10+13 running.
  • Will progressively increase the intensity in the coming days.
  • AD:
  • AD Set-up is progressing well. The low intensity from the target (-20%) seen last year has gone. Beam already on the 100MeV/c plateau (85% transmission).
  • Extraction optimization is underway.
  • n-TOF:
  • Started taking beam on 26th May (1 week late to complete installation)
  • Presently taking reduced intensity (shielding problem). 1 dedicated cycle at 5x10+12
  • Parasitic beam is set-up on the 2 available East cycles (Irradiations and North Branch)
  • SPS:
  • Beam Extracted towards LHC in both directions (TT60 & TT40) in preparation for Transfer line tests
slide6
Running with ‘Long’ FT + 3xCNGS

+ LHCPilot (When Needed)

+ MD Segment (week daytime)

slide7
Proton Sharing
  • Assuming 80% overall efficiency
  • SPS:
    • Can deliver 3.2x10+19 protons on CNGS target.
    • And ~0.5x10+19 protons to Compass (10+6 spills)
    • Beam to all other North users.
    • While delivering beam on request to LHC (several tests during the summer)
  • PS:
    • All the beams to SPS, plus …
    • 4-5 Cycles/Supercycle to Dirac – as requested.
    • Cycles to other East users according to the schedule
    • Beam to AD – every second Supercycle
    • n-TOF beam on 1 dedicated cycle + parasitic beam on all east users except Dirac. Can deliver ~7.6x10+18.
  • PS Booster:
    • All the beams to SPS and PS plus
    • Around 43% of PSB cycles to Isolde.
    • Will use frequent supercycle changes to optimize beam usage.
  • Takes account of Ion operation throughout the summer to complete the commissioning of the injectors for LHC ions.
ad